Father: When I was your age, you were 3 years old.
Son: When I reach your age, you will be 63 years old.
What is the father's present age?

Found 2 solutions by Edwin McCravy, richwmiller:
Answer by Edwin McCravy(20056)

Father: When I was your age, you were 3 years old.
Son: When I reach your age, you will be 63 years old.
Suppose the Father is x years older than the son.

Then x years ago the Father was the Son's present age, and the son was 3. 

F - x = S
S - x = 3

And x years from now the son will be the father's age and the father will
be 63.

S + x = F
F + x = 63

So we have this system:

F - x = S
S - x = 3
S + x = F
F + x = 63

The first and third equations are equivalent so we can ignore the
third equation:

F - x = S
S - x = 3
F + x = 63

Solve that system and you'll get F = 43, S = 23, x = 20

Checking: 20 years ago the father was 43-20 or 23, which is the son's present
age.  20 years ago the son was 23-20 or 3.  20 years from now the son will
be 23+20 or 43, the Father's present age, and 20 years from now the father will be 43+20 or 63.

Edwin

Answer by richwmiller(17219)

s-x=3;
f-x=s;
s+x=f;
f+x=63
f=43
s=23
x=20